The following accounting events apply to Mary’s Designs for Year 1:

Asset Source Transactions

  1. Began operations by acquiring $21,000 of cash from the issue of common stock.
  2. Performed services and collected cash of $900.
  3. Collected $4,900 of cash in advance for services to be provided over the next 12 months.
  4. Provided $12,100 of services on account.
  5. Purchased supplies of $490 on account.

Asset Exchange Transactions

  1. Purchased $4,900 of land for cash.
  2. Collected $7,900 of cash from accounts receivable.
  3. Purchased $540 of supplies with cash.
  4. Paid $3,200 for one year’s rent in advance.

Asset Use Transactions

  1. Paid $2,500 cash for salaries of employees.
  2. Paid a cash dividend of $2,200 to the stockholders.
  3. Paid off $490 of the accounts payable with cash.

Claims Exchange Transactions

  1. Placed an advertisement in the local newspaper for $180 on account.
  2. Incurred utility expense of $75 on account.

Adjusting Entries

  1. Recognized $3,675 of revenue for performing services. The collection of cash for these services occurred in a prior transaction. (See Event 3.)
  2. Recorded $1,280 of accrued salary expense at the end of Year 1.
  3. Recorded supplies expense. Had $85 of supplies on hand at the end of the accounting period.
  4. Recognized three months of expense for prepaid rent that had been used up during the accounting period.

c. Use a horizontal statements model to show how each event affects the balance sheet, income statement, and statement of cash flows. Indicate whether the event increases (+), decreases (−), or does not affect (NA) each element of the financial statements. Also, in the Statement of Cash Flows column, use the letters OA to designate operating activity, IA for investing activity, and FA for financing activity. The first event is recorded as an example.

MARY’S DESIGNS
Effect of Transactions on Financial Statements
Event Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Equity Revenue Expenses = Net Income
1. + = NA + + NA NA = NA + FA
2. = + =
3. = + =
4. = + =
5. = + =
6. = + =
7. = + =
8. = + =
9. = + =
10. = + =
11. = + =
12. = + =
13. = + =
14. = + =
15. = + =
16. = + =
17. = + =
18. = + =
Answer #1

Answer:

MARY’S DESIGNS
Effect of Transactions on Financial Statements
Event Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Equity Revenue Expenses = Net Income
1 + = NA + + NA NA = NA + FA
2 + = NA + + + NA = NA + OA
3 + = + + NA NA NA = NA + OA
4 + = NA + + + NA = NA NA NA
5 + = + + NA NA NA = NA NA NA
6 + - = NA + NA NA NA = NA - IA
7 + - = NA + NA NA NA = NA + OA
8 + - = NA + NA NA NA = NA - OA
9 - + = NA + NA NA NA = NA - OA
10 - = NA + - NA + = NA - OA
11 - = NA + - NA NA = NA - FA
12 - = - + NA NA NA = NA - OA
13 NA = + + - NA + = NA NA NA
14 NA = + + - NA + = NA NA NA
15 NA = - + + + NA = NA NA NA
16 NA = + + - NA + = NA NA NA
17 - = NA + - NA + = NA NA NA
18 - = NA + - NA + = NA NA NA

Explanations for 6,7,8,9 transactions:

Transaction Asset increase(+)/decrease(-) +/- Net affect
6 Land +, Cash - + - 0
7 Cash +, Accounts receivable - + - 0
8 Supplies +, Cash - + - 0
9 Prepaid rent +, Cash - - + 0